Write suitable functions of p and q which could be plotted against each other so as to obtain a straight line graph in the following cases.q= ap^n, q^m = ap^n, q= ab^p, q2=a(p2-b2)/p and q= AP+b+p2


1
Expert's answer
2021-07-21T09:05:55-0400

1)


"\ln q=\ln a+n\ln p"

2)


"m\ln q=\ln a+n\ln p"

3)


"\ln q=\ln a+p\ln b"